OpenStudy (christos):

integrals, can you please help me integrate sqrt(1 - 4x^2)

OpenStudy (anonymous):

can i help

OpenStudy (welshfella):

try the substitution n = 1 - 4x^2

OpenStudy (anonymous):

what are the choices

OpenStudy (christos):

I need to trigonometricly substitute

OpenStudy (welshfella):

* u = 1 - 4x^2

OpenStudy (christos):

to learn that kind of technique because I rly strucly with trigonometric substitutions

OpenStudy (christos):

strugle *

OpenStudy (christos):

can we please solve it with trigonometric substituion

OpenStudy (christos):

so that I can get the concept

OpenStudy (anonymous):

what are the choices so i can be sure thats 1 of them or do u have no choices

OpenStudy (welshfella):

yea - i'm trying to remember the appropriate substiuttion here

OpenStudy (christos):

trig substitution is the choice @youn

OpenStudy (christos):

this is not a multiple choice question, I just have to solve it @youn

OpenStudy (welshfella):

i think its x = sin theta

OpenStudy (anonymous):

thats what i got

OpenStudy (christos):

i did that but I cant get it to work @welshfella

zepdrix (zepdrix):

\[\large\rm \int\limits \sqrt{1-4x^2}~dx=\int \sqrt{1-(2x)^2}~dx\] Under the root your expression has the form: \(\large\rm 1-(stuff)^2\) Notice that if you replace the "stuff" with sin(theta), you can apply your Pythagorean Identity,\[\large\rm 1-(\sin \theta)^2=\cos^2\theta\]

OpenStudy (welshfella):

or myhbe in this case its x = sin 2theta as we have 4x^2

OpenStudy (christos):

ye but that stuff is 2x

OpenStudy (christos):

2x = sin(theta)

zepdrix (zepdrix):

Good good good :)\[\large\rm 2x=\sin \theta\]

OpenStudy (christos):

hm

OpenStudy (welshfella):

sin theta = 2x is the way to go

OpenStudy (christos):

so then its sqrt(cos^2(theta) + 2cos(theta) ?

OpenStudy (welshfella):

thats gives you sqrt ( 1 - sin^2 theta)

zepdrix (zepdrix):

\[\rm \int\limits\sqrt{1-(\color{orangered}{2x})^2}~dx\quad=\quad \int\limits\sqrt{1-(\color{orangered}{\sin \theta})^2}~dx\quad=\quad \int\limits\sqrt{\cos^2\theta}~dx\]

OpenStudy (christos):

or rather 2 * sqrt(cos(theta)) + 2cos(theta)

zepdrix (zepdrix):

Still have to deal with the `dx` from that point, but at least it makes the root easier, ya?

OpenStudy (welshfella):

so use the trig identity relating sin^2 and cos^2

OpenStudy (christos):

i dont get what I should do with dx

OpenStudy (christos):

2x = sin(theta)

OpenStudy (christos):

x = sin(theta)/2

OpenStudy (christos):

dx = cos(theta/2 right ???

OpenStudy (christos):

)*

OpenStudy (welshfella):

yea - need to find dx is terms of a trig function

zepdrix (zepdrix):

\[\large\rm dx=\frac{1}{2}\cos \theta~d \theta\]You moved the 2 to the other side? Good good good.

OpenStudy (christos):

yes and I got 2 * sqrt(cos(theta)) + 2cos(theta)

OpenStudy (christos):

is it good so far ????

zepdrix (zepdrix):

Why do you have addition?

zepdrix (zepdrix):

hmm

zepdrix (zepdrix):

\[\rm \int\limits\limits\sqrt{1-(\color{orangered}{2x})^2}~dx\quad=\quad \int\limits\limits\sqrt{1-(\color{orangered}{\sin \theta})^2}~dx\quad=\quad \int\limits\limits\sqrt{\cos^2\theta}~dx\]From that point,\[\large\rm \int\limits \cos \theta~\color{orangered}{dx}\quad=\quad \int\limits \cos \theta~\color{orangered}{\frac{1}{2}\cos \theta~d \theta}\]

zepdrix (zepdrix):

It should be a 1/2, not a 2, ya?

OpenStudy (christos):

well dx = cos(x)/2

OpenStudy (christos):

cos(theta)/2 dtheta

zepdrix (zepdrix):

Yes, you're replacing dx with... cos(theta)/2 dtheta

OpenStudy (christos):

ok so 1/2 * cos^2(x) and then half angle formula and then done

zepdrix (zepdrix):

Ummm, half angle, and a little bit of work from there. Undoing your substitution will require a triangle :) remember that step?

OpenStudy (christos):

triangle or algebra

OpenStudy (christos):

i think algebra works too

zepdrix (zepdrix):

sure sure sure

OpenStudy (welshfella):

2x = sin theta x = (1/2) sin theta dx = (1/2) cos theta

OpenStudy (christos):

thanks

OpenStudy (anonymous):

hi